OpenStudy (anonymous):

state the slope and y-intercept of the graph of each equation. y=8x+12 can someone take me step by step how to solve problems like this.

OpenStudy (anonymous):

slop 8 and intercept is 12

OpenStudy (anonymous):

thats so simple. so will x always equal the slope ?

OpenStudy (anonymous):

y=-3/5 so would the slope be -3/5 and y=y

OpenStudy (anonymous):

\[y=mx+c\] \[m\]is the slope, \[c\]is the intercept

OpenStudy (anonymous):

he's rite

OpenStudy (anonymous):

or what about 3x=-15+5y